Fermi National Laboratory


Accelerator Update for March 21, 2005 - March 23, 2005

Monday March 21
The day shift began with Operations monitoring the stack and the Recycler stash, with the Tevatron (TeV) in shot setup, with MiniBooNE taking beam, but with NuMI and Switchyard (SY) off due to shot setup.

Operations established store 4053 at 10:27 AM with an initial luminosity of 117.32E30. This store set another NEW RECORD LUMINOSITY. The Antiproton Source (Pbar) resumed stacking and the Meson MCenter experiment resumed taking beam.

Operators established beam to the Meson MTest experiment at 11:18 AM.

At 11:38 AM, MCenter experimenters reported that they had a problem with their analyzing magnet and would have to stop taking beam.

MCenter received permission at 4 PM to do straight through data runs while their analyzing magnet was off.

At 4:30 PM, Operations received a Pbar vacuum alarm, but the vacuum was good. Operators called in a controls expert.

MTest experimenters said at 5:51 PM that they were finished with the beam for the night.

Operations established beam to MCenter at 6:49 PM.

The controls expert investigating the Pbar vacuum problem identified the bad crate at 7:12 PM. Since the crate was associated with the Accumulator and the repair required a power supply adjustment, the repair will wait until after the next shot.

At 9:44 PM, the TeV suffered a sector E1 quench. Operators investigated. A TeV expert reported that a separator spark had caused the quench.

At 11:12 PM, a Booster RF station (BRF16) tripped off. It reset okay.

At 11:19 PM, BRF15 tripped off for the first off nine times in twenty minutes. After the last trip the station was bypassed.

Tuesday March 22
The midnight shift began with Pbar stacking, with the TeV turn on progressing, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:23 AM, BRF4 tripped off for the first of four times in two hours. It was left off after its last trip at 3:16 AM and BRF15 was turned on in its place.

Operations put the TeV through a dry squeeze at 1:38 AM.

At 3:27 AM, Operations reported that BRF4 & 15 were both constantly tripping off. BRF15 was the worst of the two, so operators turned it off and called in an expert. At 3:45 PM, operators began working with BRF4 to get through shot setup.

Operations established store 4056 at 4:56 AM with an initial luminosity of 78.53E30. Operations established beam to NuMI and SY at 5:02 AM. A Recycler expert began conducting reverse proton studies.

A controls expert began work on the Accumulator vacuum crate at 5:11 AM. He completed his repair and stacking resumed at 5:43 AM.

Operations reported at 7:02 AM that they were turning off Linac, Booster, the Main Injector (MI), Recycler, NuMI, MINOS, and the Meson area for accesses and maintenance work.

Operations reported at 12:08 PM that they had beam to the Linac dump.

Operations established beam to the Booster dump at 2:03 PM.

Operations established beam to MI and to MiniBooNE at 2:55 PM. Pbar resumed stacking.

At 5:36 PM, a NuMI correction magnet power supply (HV101) began tripping off. Operators investigated, but couldn't find anything obviously wrong. An expert came in and found a problem with its voltage feedback system.

Operations established beam to NuMI at 7:35 PM.

Wednesday March 23
The midnight shift began with Operations monitoring store 4056, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

At 12:48 AM, the NuMI beam permit tripped off due a drop in the HV101 current. Experts said it was okay and operators reset the permit.

At 2:28 AM, Operations reported a vacuum problem with the Recycler optical link pipe. An expert changed the oil on a vacuum pump.

At 2:30 AM, the NuMI beam permit tripped off due to another HV101 current drift. It reset okay.

Operations turned off NuMI at 7:10 AM for a scheduled access that would take all day.

At 7:20 AM, two Booster RF stations (BRF7 & 8) tripped off due to high vacuum pressure. All the vacuum pumps remained on. The stations reset okay.

The Plan for Wednesday
The plan for today is to terminate store 4056 around 9 AM and immediately go into shot setup. Pbar will stack to 40mA and then stash antiprotons into the Recycler. After the stash, Pbar experts will load some new operational curves into ARF4 and then resume stacking.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/23/2005   email Fermilab